OpenStudy (anonymous):

Find a vector equation of the line tangent to the graph of r(t) at the point Po on the curve. -> r(t)= 4costi - 3tj ; Po (2, -pi) -> r(t)= t^2i-(1/1+t)j+(4-t^2)k ; Po= (4,1,0)

OpenStudy (anonymous):

@eseidl

OpenStudy (turingtest):

take the derivative of the vector function \[\frac d{dt}\vec r(t)\]

OpenStudy (anonymous):

that i did...what value do i put in ? 2 or -pi...? @TuringTest

OpenStudy (turingtest):

find the value of t where those points occur from r(t) first

OpenStudy (turingtest):

once you know the value of t then plug that value back into r(t)

OpenStudy (turingtest):

back into r'(t)

OpenStudy (anonymous):

The x component of the point (i.e., 2) is your \[\hat{i}\]component. So,\[4\cos t=2\]in the first equation. Similarly, -pi is the j component.

OpenStudy (anonymous):

you better get the same value of t, using both:\[4\cos t=2\]and/or,\[-3t=-\pi\]

OpenStudy (anonymous):

got this part !..what about the second ? t^2=4 t= +- 2 do i put the positive value in the differentiated equation or the negative one ? @eseidl

OpenStudy (turingtest):

look at your second component to answer that

OpenStudy (anonymous):

Well, you can't tell just looking at the i component.

OpenStudy (anonymous):

@TuringTest correct.

OpenStudy (anonymous):

Which t values is consistent with\[\frac{-1}{1+t}=1\]?

OpenStudy (anonymous):

gotto go. You should have enough to answer this now.